OpenStudy (anonymous):

What is the solution of the system? y = 14x – 1 y = –11x + 4

OpenStudy (hba):

@rlawson Umm,why don't you subract them ?

OpenStudy (anonymous):

or set them equal and solve for x. Whatever method you are most comfortable with

OpenStudy (anonymous):

(5,-51) (1over5,9over5) (5over3,67over3) (3over5,37over5)

OpenStudy (hba):

@rlawson Sorry to say but you have to do the working on your own.

OpenStudy (anonymous):

set them equal, 14x - 1 = -11x + 4 and then solve for x

OpenStudy (anonymous):

x=5/3 y= 22 1/3

OpenStudy (anonymous):

thank you soo much!

OpenStudy (anonymous):

@RenaAsf do you know which answer I would pick?